OpenStudy (anonymous):

Solve the equation 1296 x = 6 Show me the steps on how to solve.

OpenStudy (anonymous):

Divide both sides by 1296; the answer is x= 1/216

OpenStudy (anonymous):

really is that it?

OpenStudy (anonymous):

Yep.

OpenStudy (amistre64):

1296 x = 6 /6 /6 ------------ 216 x = 1 /216 /216 ------------ x = 1/216

OpenStudy (anonymous):

x=just216 then?

OpenStudy (amistre64):

no, we have to follow the path that algebra takes us down x = \(\frac{1}{216}\) in the end